Grade 11General Physics

One of the sides of a rectangle is measured with error of 1% it`s area is measured with an error of 4% the error in determination of other side is

Tirupati dogga
8 Years agoGrade 11

1 Answer

Norman Manchu
8 Years ago
Area of rectangle=lxb
\therefore \frac{\Delta A}{A}=\frac{\Delta l}{l}+\frac{\Delta b}{b}
4=1+?b|b
Therefore ?b|b=3%
You are free to assume the unknown side as length or breadth too!(doesn't matter)